# What Guides Our Choices? Modeling Developers’ Trust and Behavioral Intentions Towards GenAI

### v1

@@ 4.2. Structural Model Evaluation
4.2.1. Path coefficients and significance
Table 7 presents the results of the hypotheses testing, including the mean of the bootstrap distribution (B), the standard deviation (SD), the 95% confidence interval (CI), and the p-values. The path coefficients in Fig. 1 and Table 7 are interpreted as standard regression coefficients, indicating the direct effects of one variable on another. Each hypothesis is represented by an arrow between constructs in Fig. 1. For instance, the arrow from “Functional Value” to “Trust” corresponds to H2. Given its positive path coefficient (B=0.142), genAI’s functional value is positively associated with developers’ trust in these tools. The coefficient of 0.142 indicates that when the score for functional value increases by one standard deviation unit, the score for trust increases by 0.142 standard deviation units. The analysis results (Table 7) show that most of our hypotheses are supported, except for H3 (p=0.58), H9 (p=0.06), and H10 (p=0.33). Next, we detail the factors associated with trust and behavioral intentions for the supported hypotheses with some exemplary quotes from responses to the open-ended question to illustrate our findings.
Factors associated with trust (RQ1): Our analysis supported Hypotheses H1 (p=0.00), H2 (p=0.03), and H4 (p=0.00) (Table 7). First, the support for system/output quality in fostering trust (H1) can be explained by how developers prefer tools that deliver accurate, reliable outputs matching their work style and practices (Wang et al., 2023; Yu et al., 2019). Next, the functional value of genAI, encompassing educational benefits and practical advantages, promotes trust (H2) since developers prioritize tools that offer tangible utility in their work (Johnson et al., 2023; Ziegler et al., 2024). For instance, one survey respondent mentioned the practical utility of genAI tools, stating, “I find value in these models for creative endeavors, gaining different perspectives, or coming up with ideas I wouldn’t have otherwise”. Finally, goal maintenance is relevant for cultivating trust (H4). The alignment between a developer’s goals and genAI’s actions supports using genAI tools to achieve these goals. This eliminates the need for developers to constantly verify the relevance of genAI’s outputs, thereby reducing cognitive load. This congruence ultimately enhances genAI’s credibility as a cognitive collaborator (Wischnewski et al., 2023) rather than as an independent and potentially untrustworthy tool, thus bolstering trust in these tools.
@@ 4.2.1. Path coefficients and significance
Trust (H5) is pivotal in shaping adoption decisions as it reduces resistance to new technologies (Xiao et al., 2014; Witschey et al., 2015). When developers trust genAI tools, they perceive them as credible partners, enhancing their willingness to use these tools. Moreover, developers’ cognitive styles significantly shape their intentions to adopt genAI tools. Developers motivated by the intrinsic enjoyment of technology (H6) have higher intentions to adopt genAI tools. In contrast, those with a task-oriented approach tend to be more cautious and hesitant about the cognitive effort they are willing to invest in these tools (Burnett et al., 2016). One respondent echoed this, stating, “I am slow to adopt new workflows; I put off actively exploring new tools unless it is related to what I need to do”. Higher computer self-efficacy within peer groups is also significantly associated with increased intentions to use genAI tools (H7). Despite generally high self-efficacy, some developers face interaction challenges with genAI that may impact their confidence and adoption rates. One respondent shared, “I see my colleagues getting good responses while I fiddle around to get the answers I need. Also, it does not always show the right document relevant to me, so I prefer traditional ways”. Furthermore, we found that developers with higher risk tolerance are significantly more inclined to use these tools than risk-averse individuals (H8). The context (and involved stakes) in which these tools are used further play a role, as highlighted by another respondent: “I don’t use it yet to write code that I can put my name behind in production; I just use it for side projects or little scripts to speed up my job, but not in actual production code”.
Finally, our analysis supported Hypothesis H11 (p=0.00), highlighting a significant positive association between developers’ behavioral intention to use genAI tools and its usage in their work. This corroborates with prior technology acceptance models (Venkatesh et al., 2003, 2012), emphasizing the pivotal role of behavioral intentions in predicting use behavior.
Control variables: Although experience is often relevant for technology adoption (Venkatesh et al., 2003), our analysis found no significant associations between SE experience and trust, behavioral intentions, or usage of genAI tools. This is likely since genAI introduces a new interaction paradigm (Weisz et al., 2023), which diverges from traditional SE tools and requires different skills and interactions not necessarily linked to SE experience. Familiarity with genAI, while potentially influential, was excluded as a control variable due to a highly skewed distribution of responses, with most participants reporting high familiarity. Including such skewed variables could lead to unreliable estimates and compromise the model’s validity (Hair et al., 2019; Sarstedt and Cheah, 2019). Similarly, the gender variable was excluded due to its skewed distribution. The analysis of unobserved heterogeneity (see supplemental (sup, [n. d.])) confirms the absence of any group differences in the model (e.g., organizational heterogeneity) caused by unmeasured criteria.
Table 7. Standarized path coefficients (B), standard deviations (SD), confidence intervals (CI), p values, and effect sizes ($f^{2}$)
@@ B SD 95% CI p $f^{2}$
  • BI: Behavioral Intention. We consider $f^{2}<$ 0.02 to be no effect, $f^{2}\in$ [0.02, 0.15) to be small, $f^{2}\in$ [0.15, 0.35) to be medium, and $f^{2}>$ 0.35 to be large (Cohen, 2013).
4.2.2. Model evaluation
@@
We assessed the relationship between constructs and the predictive capabilities of the theoretical model by evaluating the model’s explanatory power ($R^{2}$, Adjusted (Adj.) $R^{2}$), model fit (SRMR), effect sizes ($f^{2}$), and predictive relevance ($Q^{2}$) (Russo and Stol, 2021).
@@
Explanatory power: The coefficient of determination ($R^{2}$ and Adj. $R^{2}$ values) indicate the proportion of variance in the endogenous variables explained by the predictors. Ranging from 0 to 1, higher $R^{2}$ values signify greater explanatory power, with 0.25, 0.5, and 0.75 representing weak, moderate, and substantial levels, respectively (Hair et al., 2019). As shown in Table 8, the $R^{2}$ values in our model are 0.68 for Trust, 0.66 for Behavioral intention, and 0.33 for Usage, demonstrating moderate to substantial explanatory power, well above the accepted threshold of 0.19 (Chin et al., 1998). Further, Table 7 presents the effect sizes ($f^{2}$), which measure the impact of each predictor on the endogenous variables. The effect sizes indicate that the predictors exhibit medium to large effects on their respective endogenous variables for all supported hypotheses in our model, with values ranging from 0.14 to 0.54 (Cohen, 2013), further corroborating the model’s explanatory power. [^3] Large $R^{2}$ and $f^{2}$ can occasionally indicate overfitting. We thoroughly evaluated this issue by analyzing residuals and conducting cross-validation, finding no evidence of model overfitting (see (sup, [n. d.])).
@@
Model fit: We analyzed the overall model fit using the standardized root mean square residual (SRMR), a recommended fit measure for detecting misspecification in PLS-SEM models (Russo and Stol, 2021). Our results suggest a good fit of the data in the theoretical model, with SRMR = 0.077, which is below the suggested thresholds of 0.08 (conservative) and 0.10 (lenient) (Henseler et al., 2016).
@@
Predictive relevance: Finally, we evaluated the model’s predictive relevance using Stone-Geisser’s $Q^{2}$ (Stone, 1974), a measure of external validity (Hair et al., 2019) obtainable via the PLS-predict algorithm (Shmueli et al., 2016) in SmartPLS. PLS-predict is a holdout sample-based procedure: it divides the data into $k$ subgroups (folds) of roughly equal size, using (k-1) folds as a training sample to estimate the model, while the remaining fold serves as a holdout to assess out-of-sample predictive power. $Q^{2}_{\text{predict}}$ values are calculated for endogenous variables; values greater than 0 indicate predictive relevance, while negative values suggest the model does not outperform a simple average of the endogenous variable. Our sample was segmented into k=10 parts, and 10 repetitions were used to derive the $Q^{2}_{\text{predict}}$ statistic (Hair et al., 2019), all of which were greater than 0 (Table 8), confirming our model’s adequacy in terms of predictive relevance.
Table 8. Coefficient of determination and predictive relevance
@@
4.2.3. Common method bias
We collected data via a single survey instrument, which might raise concerns about Common Method Bias/Variance (CMB/CMV) (Russo and Stol, 2021). To test for CMB, we applied Harman’s single factor test (Podsakoff et al., 2003) on the latent variables. No single factor explained more than 23% variance. An unrotated exploratory factor analysis with a forced single-factor solution was conducted, which explained 30.3% of the variance, well below the 50% threshold. Additionally, we used Kock’s collinearity approach (Kock, 2015). The VIFs for the latent variables ranged from 1.01 to 2.45, all under the cut-off of 3.3. These indicate that CMB was not a concern in our study.
5. Discussion
5.1. Implications for practice
@@
To achieve goal maintenance, the AI should consistently account for the developer’s (1) current state; (2) immediate goals, and the expected success (outcomes from AI) (Wischnewski et al., 2023); as well as (3) preferences for transitioning from their current state (1) to achieving their immediate goals (2). These preferences may involve process methodologies, interaction styles, output specifications, alignment with the work style, or safety/security considerations (Johnson et al., 2023; Burnett et al., 2016). Given that an individual’s goal(s) may evolve in response to changes in their task state, toolsmiths must also prioritize adaptability in genAI tool design. For instance, adaptability can be incorporated by iteratively adjusting AI’s actions based on user input on (1), (2), and (3) to ensure its ongoing alignment with their shifting goals.
Allowing developers the flexibility to explicitly steer AI’s actions as needed is also important for goal maintenance. This control can be essential if the genAI tool deviates from the expected trajectory, enabling developers to (re)calibrate it to support their goals. Further, this can also support developers’ metacognitive flexibility (Tankelevitch et al., 2024), i.e., they may adapt their cognitive strategies based on new information or task-state changes. The ability to steer the AI to align with these adapted strategies helps accommodate their evolving goals.
Design for contextual transparency. Developers often face decisions about incorporating genAI tool support for their tasks. Any mismatch between their expectations and genAI’s true capabilities can lead to over/under-estimating the tool’s functionality, increasing the risk of errors, lost productivity, or potential adverse outcomes in critical tasks (Pearce et al., 2022). Such outcomes are detrimental to fostering subsequent trust. Thus, helping developers to calibrate their expectations to match the true tool capabilities is essential. Creating this alignment involves designing interfaces that explicitly communicate the tool’s capabilities and limitations, consistent with HAI interaction guidelines (Amershi et al., 2019; Google, 2023). This clarity allows individuals to form accurate expectations about system quality (H1) within their task contexts and assess its functional value (H2), thereby fostering appropriate trust (Jacovi et al., 2021). To that end, we suggest:
@@
(b) Communicating genAI’s limitations for the current task and scoping assistance under conditions of uncertainty to incite warranted trust (Jacovi et al., 2021). Doing so allows developers to consider the tool’s practical utility for their task context, clarifying its appropriate functional value (H2) in their work.
Inclusive tool design for HAI-UX fairness. AI fairness has gained substantial traction over the years (Green and Chen, 2019). While much of the research and discussion has focused on data or algorithmic fairness, fairness should also include user experiences in human-AI interactions (HAI-UX). We advocate promoting fairness in HAI-UX through inclusive genAI tool design, specifically by supporting developers’ diverse cognitive styles.
Our findings indicate that developers who are motivated to use genAI tools for their own sake (H6), have higher computer self-efficacy (H7), and have greater tolerance for risk (H8) are more inclined to adopt these tools compared to their peers. To achieve more equitable acceptance of these tools across the cognitive spectrum of developers, future designs must prioritize adaptability based not only on developers’ goals (discussed above) but also on their cognitive styles. This can be achieved by capturing these styles (e.g., using the survey questions from this study) and designing genAI tools that dynamically adapt to align with these styles using various strategies. For example, to support developers motivated by task completion, genAI tools can solicit their immediate goals and expected outcomes and deliver contextually appropriate information (and explanations) consistent with their preferred styles. This would simplify developers’ ongoing tasks without overwhelming them with unrelated features or extraneous information, helping them complete their tasks effectively. As another example, genAI tools can support individuals with lower self-efficacy by offering explicit cues that differentiate between errors arising from prompting issues and those due to system limitations. For instance, it can caution users about tasks where it typically underperforms, based on prior feedback, or highlight specific parts of the prompt that influenced the generated output. This distinction can help prevent individuals from doubting their ability to effectively use these tools in their work.
5.2. Implications for research
@@
Non-significant associations: Our analysis did not find support for Hypotheses H3 (p=0.59), H9 (p=0.06), and H10 (p=0.33). These findings are surprising, as ease of use, information processing, and tinkering learning style are relevant when considering traditional software tools (Burnett et al., 2016; Venkatesh et al., 2003). However, in genAI contexts, these constructs may manifest differently due to the altered dynamics of user engagement compared to more traditional software. The intuitive nature of genAI interfaces might diminish the traditional impact of these factors. For example, ease of use might not show a relation as using these interfaces is inherently easy; instead, the appropriateness of the queries is what matters. Similarly, developers’ information processing style (H9) did not significantly influence their intentions to use genAI tools, likely because how individuals articulate their needs—a single comprehensive prompt or sequence of queries—often aligns with their preferences for consuming information (comprehensive or selective). The lack of a relationship for tinkering style (H10), as well, could be attributed to genAI’s interaction paradigm, which is primarily centered around (re)formulating and following up with queries rather than “tinkering” with the software’s features. If these speculations hold, how certain validated constructs were framed in the current study (sup, [n. d.]) might have indeed limited our understanding of these dynamics. Future research should explore these constructs more deeply within the context of human-genAI interactions. For instance, instead of focusing on ‘ease of use’ or ‘tinkering with software features’, studies could examine ‘ease of prompting’ or ‘tinkering with prompt strategies’ and how preferences (and proficiency) in these areas influence developers’ trust and behavioral intentions. Understanding these dynamics can inform the future design and adoption strategies of genAI tools, aligning them more closely with user interaction patterns and cognitive styles.
5.3. Threats to validity and limitations

### v2

@@ IV-B Structural Model Evaluation
IV-B1 Path coefficients and significance
Table VII presents the results of the hypotheses testing, including the mean of the bootstrap distribution (B), the standard deviation (SD), the 95% confidence interval (CI), and the p-values. The path coefficients in Fig. 1 and Table VII are interpreted as standard regression coefficients, indicating the direct effects of one variable on another. Each hypothesis is represented by an arrow between constructs in Fig. 1. For instance, the arrow from “Functional Value” to “Trust” corresponds to H2. Given its positive path coefficient (B=0.142), genAI’s functional value is positively associated with developers’ trust in these tools. The coefficient of 0.142 indicates that when the score for functional value increases by one standard deviation unit, the score for trust increases by 0.142 standard deviation units. The analysis results (Table VII) show that most of our hypotheses are supported, except for H3 (p=0.58), H9 (p=0.06), and H10 (p=0.33). Next, we detail the factors associated with trust and behavioral intentions for the supported hypotheses with some exemplary quotes from responses to the open-ended question to illustrate our findings.
Factors associated with trust (RQ1): Our analysis supported Hypotheses H1 (p=0.00), H2 (p=0.03), and H4 (p=0.00) (Table VII). First, the support for system/output quality in fostering trust (H1) can be explained by how developers prefer tools that deliver accurate, reliable outputs matching their work style and practices [14, 67]. Next, the functional value of genAI, encompassing educational benefits and practical advantages, promotes trust (H2) since developers prioritize tools that offer tangible utility in their work [16, 69]. For instance, one survey respondent mentioned the practical utility of genAI tools, stating, “I find value in these models for creative endeavors, gaining different perspectives, or coming up with ideas I wouldn’t have otherwise”. Finally, goal maintenance is relevant for cultivating trust (H4). The alignment between a developer’s goals and genAI’s actions supports using genAI tools to achieve these goals. This eliminates the need for developers to constantly verify the relevance of genAI’s outputs, thereby reducing cognitive load. This congruence ultimately enhances genAI’s credibility as a cognitive collaborator [34] rather than as an independent and potentially untrustworthy tool, thus bolstering trust in these tools.
@@ IV-B1 Path coefficients and significance
Trust (H5) is pivotal in shaping adoption decisions as it reduces resistance to new technologies [74, 75]. When developers trust genAI tools, they perceive them as credible partners, enhancing their willingness to use these tools. Moreover, developers’ cognitive styles significantly shape their intentions to adopt genAI tools. Developers motivated by the intrinsic enjoyment of technology (H6) have higher intentions to adopt genAI tools. In contrast, those with a task-oriented approach tend to be more cautious and hesitant about the cognitive effort they are willing to invest in these tools [26]. One respondent echoed this, stating, “I am slow to adopt new workflows; I put off actively exploring new tools unless it is related to what I need to do”. Higher computer self-efficacy within peer groups is also significantly associated with increased intentions to use genAI tools (H7). Despite generally high self-efficacy, some developers face interaction challenges with genAI that may impact their confidence and adoption rates. One respondent shared, “I see my colleagues getting good responses while I fiddle around to get the answers I need. Also, it does not always show the right document relevant to me, so I prefer traditional ways”. Furthermore, we found that developers with higher risk tolerance are significantly more inclined to use these tools than risk-averse individuals (H8). The context (and involved stakes) in which these tools are used further play a role, as highlighted by another respondent: “I don’t use it yet to write code that I can put my name behind in production; I just use it for side projects or little scripts to speed up my job, but not in actual production code”.
Finally, our analysis supported Hypothesis H11 (p=0.00), highlighting a significant positive association between developers’ behavioral intention to use genAI tools and its usage in their work. This corroborates with prior technology acceptance models [21, 22], emphasizing the pivotal role of behavioral intentions in predicting use behavior.
TABLE VII: Standarized path coefficients (B), standard deviations (SD), confidence intervals (CI), p values, and effect sizes ($f^{2}$)
@@ B SD 95% CI p $f^{2}$
  • BI: Behavioral Intention. We consider $f^{2}<$ 0.02 to be no effect, $f^{2}\in$ [0.02, 0.15) to be small, $f^{2}\in$ [0.15, 0.35) to be medium, and $f^{2}>$ 0.35 to be large [90].
Control variables: Although experience is often relevant for technology adoption [21], our analysis found no significant associations between SE experience and trust, behavioral intentions, or usage of genAI tools. This is likely since genAI introduces a new interaction paradigm [44], which diverges from traditional SE tools and requires different skills and interactions not necessarily linked to SE experience. Familiarity with genAI, while potentially influential, was excluded as a control variable due to a highly skewed distribution of responses, with most participants reporting high familiarity. Including such skewed variables could lead to unreliable estimates and compromise the model’s validity [52, 91]. Similarly, the gender variable was excluded due to its skewed distribution. The analysis of unobserved heterogeneity (see supplemental [53]) confirms the absence of any group differences in the model (e.g., organizational heterogeneity) caused by unmeasured criteria.
IV-B2 Model evaluation
@@
We assessed the relationship between constructs and the predictive capabilities of the theoretical model by evaluating the model’s explanatory power ($R^{2}$, Adjusted (Adj.) $R^{2}$), model fit (SRMR), effect sizes ($f^{2}$), and predictive relevance ($Q^{2}$) [85].
@@
Explanatory power: The coefficient of determination ($R^{2}$ and Adj. $R^{2}$ values) indicate the proportion of variance in the endogenous variables explained by the predictors. Ranging from 0 to 1, higher $R^{2}$ values signify greater explanatory power, with 0.25, 0.5, and 0.75 representing weak, moderate, and substantial levels, respectively [52]. As shown in Table VIII, the $R^{2}$ values in our model are 0.68 for Trust, 0.66 for Behavioral intention, and 0.33 for Usage, demonstrating moderate to substantial explanatory power, well above the accepted threshold of 0.19 [92]. Further, Table VII presents the effect sizes ($f^{2}$), which measure the impact of each predictor on the endogenous variables. The effect sizes indicate that the predictors exhibit medium to large effects on their respective endogenous variables for all supported hypotheses in our model, with values ranging from 0.14 to 0.54 [90], further corroborating the model’s explanatory power. [^3] Large $R^{2}$ and $f^{2}$ can occasionally indicate overfitting. We thoroughly evaluated this issue by analyzing residuals and conducting cross-validation, finding no evidence of model overfitting (see [53]).
@@
Model fit: We analyzed the overall model fit using the standardized root mean square residual (SRMR), a recommended fit measure for detecting misspecification in PLS-SEM models [85]. Our results suggest a good fit of the data in the theoretical model, with SRMR = 0.077, which is below the suggested thresholds of 0.08 (conservative) and 0.10 (lenient) [93].
@@
Predictive relevance: Finally, we evaluated the model’s predictive relevance using Stone-Geisser’s $Q^{2}$ [94], a measure of external validity [52] obtainable via the PLS-predict algorithm [95] in SmartPLS. PLS-predict is a holdout sample-based procedure: it divides the data into $k$ subgroups (folds) of roughly equal size, using ($k-1$) folds as a training sample to estimate the model, while the remaining fold serves as a holdout to assess out-of-sample predictive power. $Q^{2}_{\text{predict}}$ values are calculated for endogenous variables; values greater than 0 indicate predictive relevance, while negative values suggest the model does not outperform a simple average of the endogenous variable. Our sample was segmented into k=10 parts, and 10 repetitions were used to derive the $Q^{2}_{\text{predict}}$ statistic [52], all of which were greater than 0 (Table VIII), confirming our model’s adequacy in terms of predictive relevance.
TABLE VIII: Coefficient of determination and predictive relevance
@@
IV-B3 Common method bias
We collected data via a single survey instrument, which might raise concerns about Common Method Bias/Variance (CMB/CMV) [85]. To test for CMB, we applied Harman’s single factor test [96] on the latent variables. No single factor explained more than 23% variance. An unrotated exploratory factor analysis with a forced single-factor solution was conducted, which explained 30.3% of the variance, well below the 50% threshold. Additionally, we used Kock’s collinearity approach [97]. The VIFs for the latent variables ranged from 1.01 to 2.45, all under the cut-off of 3.3. These indicate that CMB was not a concern in our study.
V Discussion
V-A Implications for practice
@@
To achieve goal maintenance, the AI should consistently account for the developer’s (1) current state; (2) immediate goals, and the expected success (outcomes from AI) [34]; as well as (3) preferences for transitioning from their current state (1) to achieving their immediate goals (2). These preferences may involve process methodologies, interaction styles, output specifications, alignment with the work style, or safety/security considerations [16, 26]. Given that an individual’s goal(s) may evolve in response to changes in their task state, toolsmiths must also prioritize adaptability in genAI tool design. For instance, adaptability can be incorporated by iteratively adjusting AI’s actions based on user input on (1), (2), and (3) to ensure its ongoing alignment with their shifting goals.
Allowing developers the flexibility to explicitly steer AI’s actions as needed is also important for goal maintenance. This control can be essential if the genAI tool deviates from the expected trajectory, enabling developers to (re)calibrate it to support their goals. Further, this can also support developers’ metacognitive flexibility [98], i.e., they may adapt their cognitive strategies based on new information or task-state changes. The ability to steer the AI to align with these adapted strategies helps accommodate their evolving goals.
Design for contextual transparency. Developers often face decisions about incorporating genAI tool support for their tasks. Any mismatch between their expectations and genAI’s true capabilities can lead to over/under-estimating the tool’s functionality, increasing the risk of errors, lost productivity, or potential adverse outcomes in critical tasks [12]. Such outcomes are detrimental to fostering subsequent trust. Thus, helping developers to calibrate their expectations to match the true tool capabilities is essential. Creating this alignment involves designing interfaces that explicitly communicate the tool’s capabilities and limitations, consistent with HAI interaction guidelines [99, 100]. This clarity allows individuals to form accurate expectations about system quality (H1) within their task contexts and assess its functional value (H2), thereby fostering appropriate trust [35]. To that end, we suggest:
@@
(b) Communicating genAI’s limitations for the current task and scoping assistance under conditions of uncertainty to incite warranted trust [35]. Doing so allows developers to consider the tool’s practical utility for their task context, clarifying its appropriate functional value (H2) in their work.
Inclusive tool design for HAI-UX fairness. AI fairness has gained substantial traction over the years [101]. While much of the research and discussion has focused on data or algorithmic fairness, fairness should also include user experiences in human-AI interactions (HAI-UX). We advocate promoting fairness in HAI-UX through inclusive genAI tool design, specifically by supporting developers’ diverse cognitive styles.
Our findings indicate that developers who are motivated to use genAI tools for their own sake (H6), have higher computer self-efficacy (H7), and have greater tolerance for risk (H8) are more inclined to adopt these tools compared to their peers. To achieve more equitable acceptance of these tools across the cognitive spectrum of developers, future designs must prioritize adaptability based not only on developers’ goals (discussed above) but also on their cognitive styles. This can be achieved by capturing these styles (e.g., using the survey questions from this study) and designing genAI tools that dynamically adapt to align with these styles using various strategies. For example, to support developers motivated by task completion, genAI tools can solicit their immediate goals and expected outcomes and deliver contextually appropriate information (and explanations) consistent with their preferred styles. This would simplify developers’ ongoing tasks without overwhelming them with unrelated features or extraneous information, helping them complete their tasks effectively. As another example, genAI tools can support individuals with lower self-efficacy by offering explicit cues that differentiate between errors arising from prompting issues and those due to system limitations. For instance, it can caution users about tasks where it typically underperforms, based on prior feedback, or highlight specific parts of the prompt that influenced the generated output. This distinction can help prevent individuals from doubting their ability to effectively use these tools in their work. Further, for risk-averse developers, transparency in genAI behavior and outputs should be emphasized. GenAI should also provide explicit indicators of any potential flaws or uncertainties in its outputs (e.g., confidence scores [14]). Such transparency can help developers make informed decisions about incorporating these outputs into their work (e.g., using AI-generated code), accommodating the caution and deliberation associated with risk aversion.
V-B Implications for research
@@
Non-significant associations: Our analysis did not find support for Hypotheses H3 (p=0.59), H9 (p=0.06), and H10 (p=0.33). These findings are surprising, as ease of use, information processing, and tinkering learning style are relevant when considering traditional software tools [26, 21]. However, in genAI contexts, these constructs may manifest differently due to the altered dynamics of user engagement compared to more traditional software. The intuitive nature of genAI interfaces might diminish the traditional impact of these factors. For example, ease of use might not show a relation as using these interfaces is inherently easy; instead, the appropriateness of the queries is what matters. Similarly, developers’ information processing style (H9) did not significantly influence their intentions to use genAI tools, likely because how individuals articulate their needs—a single comprehensive prompt or sequence of queries—often aligns with their preferences for consuming information (comprehensive or selective). The lack of a relationship for tinkering style (H10), as well, could be attributed to genAI’s interaction paradigm, which is primarily centered around (re)formulating and following up with queries rather than “tinkering” with the software’s features. If these speculations hold, how certain validated constructs were framed in the current study [53] might have indeed limited our understanding of these dynamics. Future research should explore these constructs more deeply within the context of human-genAI interactions. For instance, instead of focusing on ‘ease of use’ or ‘tinkering with software features’, studies could examine ‘ease of prompting’ or ‘tinkering with prompt strategies’ and how preferences (and proficiency) in these areas influence developers’ trust and behavioral intentions. Understanding these dynamics can inform the future design and adoption strategies of genAI tools, aligning them more closely with user interaction patterns and cognitive styles.
V-C Threats to validity and limitations